OpenStudy (anonymous):

Simplify the expression... 4/((3x)^-2)

jhonyy9 (jhonyy9):

4/1/9x^2 = 36x^2

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

Sorry, got it. Thank you :)

jhonyy9 (jhonyy9):

yw good luck bye - so i am andrew too

OpenStudy (anonymous):

lol xD

OpenStudy (anonymous):

\[4/((3x){}^{\wedge}-2)=4\left/\left(\frac{1}{9 x^2}\right)\right.=36 x^2 \]